# Towards Human-AI Collaboration in Healthcare: Guided Deferral Systems with Large Language Models

### v1

@@ 4.2 Pilot Study Results
Figure 2: Human accuracy without guidance is lower and more variable than with guidance (HAIC). Accuracy of prediction methods from pilot study. Guided humans outperform both unguided humans and the LLM.
@@ 5 Conclusion
We develop a guided deferral framework using LLMs for clinical decisions, with HAIC as a key focus. Our study shows that instruction-tuned small-scale LLMs can achieve significant deferral and classification performance while maintaining computational efficiency through instruction-tuning on supervised data from a large-scale LLM. We utilise two prediction sources—verbalised and hidden-state, yielding valuable insights for classification and deferral. Finally, we validate the efficacy of our proposed guided deferral system through a pilot study, with the results showing a significant improvement in human decision-making performance under the LLM guidance.
Limitations.
In this paper, we prioritise healthcare in our analysis due to its greater ethical significance in research. However, there exists a severe lack of high-quality report data with labels in medical domain, making it difficult for LLM evaluation. When implementing our system, extra efforts are required to educate clinicians about its capabilities to prevent unexpected errors. This involves addressing human biases such as anchoring and confirmation bias, as well as briefing them on the system’s training data distribution, including concepts such as data drift and out-of-distribution instances. Such understanding is vital, particularly in scenarios where variations in report style or format might affect the system’s performance.
Acknowledgements.
The OSCLMRIC dataset was collected by Professor Jeremy Fairbank with HRA approval (IRAS Project ID: 207858, Protocol Number:12139). The authors extend their gratitude to Professor. Andrew Zisserman and Dr. Amir Jamaludin from the Visual Geometry Group, University of Oxford, for the provision of data. JS is supported by the EPSRC Center for Doctoral Training in Health Data Science (EP/S02428X/1). JS gratefully acknowledges the Kellogg College Research Support Grant of which supported this research. The authors acknowledge UKRI grant reference EP/X040186/1 (Turing AI Fellowship: Ultra Sound Multi-Modal Video-based Human-Machine Collaboration).
@@ Appendix A Additional Listings
Listing 2: Base-prompt used in generating data to instruction-tune and in parsing of reports. The relevant level, disorder, question, and report were inserted to make the final prompt.
@@ Conclusion:
Listing 3: An example spinal MRI radiological report.
@@ D.1 Hidden-State Classifier Experiments
D.2 Verbalised and Hidden-State Predictions Contribute Distinct Pieces of Information to Classification.
In analysing the individual effects $\hat{t}$ and $\hat{\epsilon}$, we focus our analysis on the cases when these predictions differ when their combined prediction $\hat{\mu}$ is correct and incorrect. Initially, we find a positive correlation between the two predictions: 0.53 ± 0.07.
$\hat{\epsilon}$ corrects predictions of $\hat{t}$.
We find that, when $\hat{\mu}$ is correct and the predictions of $\hat{t}$ and $\hat{\epsilon}$ differ, a correct $\hat{\epsilon}$ prediction is combining with an incorrect $\hat{t}$ prediction in the majority of cases, at an average of (75 ± 0.14)%. When $\hat{\mu}$ is incorrect, a correct $\hat{t}$ prediction is combining with an incorrect $\hat{\epsilon}$ prediction in the majority of cases, at an average of (75 ± 0.25)%. See Figure LABEL:fig:mu_correct_apd for confusion matrices.
When inspecting the generated textual contents, we see that in the cases where $\hat{\epsilon}$ is correct and $\hat{t}$ is incorrect, it is largely due to the LLM either violating an assumption given in the prompt. Regardless, the logic in the reasoning for the verbalised prediction is always sound.

### v2

@@ 4.2 Pilot Study Results
Figure 4: Human accuracy without guidance is lower and more variable than with guidance (HAIC). Accuracy of prediction methods from pilot study. On average, guided humans outperform both unguided humans and the LLM.
@@ 5 Conclusion
We develop a guided deferral framework using LLMs for clinical decisions, with HAIC as a key focus. Our study shows that IT small-scale LLMs can achieve significant deferral and classification performance while maintaining computational efficiency through instruction-tuning on supervised data from a large-scale LLM. We utilise two prediction sources—verbalised and hidden-state, yielding valuable insights for classification and deferral. The combination of these sources results in a significant classification and high-uncertainty deferral prediction. Finally, we validate the efficacy of our proposed guided deferral system through a pilot study, with the results showing a significant improvement in human decision-making performance under the LLM guidance.
Limitations. In this paper, we prioritise healthcare in our analysis due to its greater ethical significance in research. However, there exists a severe lack of high-quality report data with labels in medical domain, making it difficult for LLM evaluation. When implementing our system, extra efforts are required to educate clinicians about its capabilities to prevent unexpected errors. This involves addressing human biases such as anchoring and confirmation bias, as well as briefing them on the system’s training data distribution, including concepts such as data drift and out-of-distribution instances. Such understanding is vital, particularly in scenarios where variations in report style or format might affect the system’s performance.
Acknowledgements. JS is supported by the EPSRC Center for Doctoral Training in Health Data Science (EP/S02428X/1). JS gratefully acknowledges the Kellogg College Research Support Grant of which supported this research. The authors acknowledge UKRI grant reference EP/X040186/1 (Turing AI Fellowship: Ultra Sound Multi-Modal Video-based Human-Machine Collaboration). The OSCLMRIC dataset was collected by Professor Jeremy Fairbank with HRA approval (IRAS Project ID: 207858, Protocol Number:12139). The authors extend their gratitude to Professor. Andrew Zisserman and Dr. Amir Jamaludin from the Visual Geometry Group, University of Oxford, for the provision of this data. A special thank-you to all participants of the pilot study.
@@ Appendix A Additional Listings
Listing 1: Base-prompt used in generating data to instruction-tune and in parsing of reports. The relevant level, disorder, question, and report were inserted to make the final prompt.
@@ Conclusion:
Listing 2: An example spinal MRI radiological report.